Solve
x/3 - 1/4= 3x-8/12

Thank You!!!

Answer by Nate(3500)

x/3 - 1/4 = 3x - 8/12
12x/3 - 12/4 = 3(12)x - 8(12)/12
4x - 3 = 36x - 8
5 = 32x
5/32 = x = 0.15625